OpenStudy (anonymous):

Rodrigo is looking at the system below. 2x + 3y = 9 x - 8y = 11 To solve, he first multiplied the second equation by two and then subtracted the equations. Is Rodrigo correct? Justify your answer. If incorrect, describe the error.

OpenStudy (anonymous):

@Dbzfan836

OpenStudy (anonymous):

i hate these

OpenStudy (dbzfan836):

use geogebra, a free graphing software. enter things like 2x+3y=9, and it will help alot

OpenStudy (anonymous):

ive acyually tried that but it nearly destroyed my pc

OpenStudy (anonymous):

idk how but it did

OpenStudy (dbzfan836):

no, thats a fake one. heres the link: ( ͡° ͜ʖ ͡°) www.geogebra.org/cms/en/download

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

how much are you typing @texaschic101 ??

OpenStudy (texaschic101):

you can multiply the 2nd equation by -2, and then add them... 2x + 3y = 9 x - 8y = 11---->(-2) ---------------- 2x + 3y = 9 -2x + 16y = -22 ---------------add 19y = - 13 y = -13/19 OR you can multiply second equation by 2, then subtract them 2x + 3y = 9 x - 8y = 11---->(2) ---------------- 2x + 3y = 9 2x - 16y = 22 --------------subtract 2x + 3y = 9 -2x + 16y = -22 -------------- 19y = -13 y = -13/19 yes....he is correct if y = -13/19, sub that in and solve for x 2x + 3y = 9 2(-13/19) + 3y = 9 -26/19 + 3y = 9 (multiply by 19) -26 + 57y = 171 57y = 171 + 26 57y = 197 y = 197/57 check.. 2x + 3y = 9 2(-13/19) + 3(197/57) = 9 -26/19 + 591/57 = 9 -78/57 + 591/57 = 9 513/57 = 9 9 = 9 (correct) so x = -13/19 and y = 197/57 and Rodrigo is correct

OpenStudy (dbzfan836):

oh ok

OpenStudy (anonymous):

oh flutter

OpenStudy (anonymous):

you know i didnt mean flutter

OpenStudy (texaschic101):

lol

OpenStudy (anonymous):

but thank you very much miss texas

OpenStudy (texaschic101):

sorry it took so long...not the fastest at subtracting them....I do better with adding them

OpenStudy (anonymous):

its completely ok